OpenStudy (anonymous):

What is the oxidation number of phosphorus (P) in phosphoric acid (H3PO4)?

OpenStudy (anonymous):

if O is -2 and hydrogen +1 then phosphorus is +5

OpenStudy (anonymous):

For acids the naming follows the oxidation state of the central atom (phosphorus in this case). Below is a list of the oxidation numbers and their corresponding prefix/suffix combination: +1, +2: hypo_______ous +3, +4: _______ous +5, +6: _______ic +7: per________ic What this tells you is that the oxidation number of P in phosphorus acid must be +3 or +4, while the value for phosphoric acid must be +5 or +6. Since phosphorus is a member of group 5A, it has 5 electrons in its valence shell. As a result, the most common oxidation states it can have are +3 (s2p0) or +5 (s0p0). So far we have determined the oxidation state at P. What needs to be done now is to add as many oxygens and hydroxyl (OH) to make the molecule neutral. The correct combination will have the correct Lewis-dot structure. For phosphorous acid we need a combination that will add up to +3. This can be done by adding 3 OH- to the central atom to yield the structure H3PO3. There is a little caveat though. Because this is not a hydro___acid, it is implied that there must be at least one oxo ligand (O^2-) bonded to P. Therefore, the actual bonding structure is not P(OH)3 but rather H-P(=O)(OH)2, where one H is bonded directly to the phosphorus atom and is the least acidic of the protons. The great thing is that the oxidation charge of P is still +3 because P is slightly more electronegative than H (some theories will say otherwise); thus the hydrogen is regarded as H+ for this example. With phosphoric acid, charge of +5, you can have 3 OH- and 1 O^2- to make a neutral molecule: O=P(OH)3. I hope this sort of helped you out. These molecules are certainly the least straightforward to apply this rule to, but it is good to think about it :)

OpenStudy (anonymous):

Or did you want this?. H is always +1 so the H's have a +3 charge O is always -2 so the O's have a -8 charge that gives the compound a total charge of -5 without the chrage of P yet. To have the entir ecompound have a -1 charge the oxidation number for P is +4. *except H3PO4 is phosphoric acid and the charge is zero so if that is the compound youre asking for the oxidation number is +5

OpenStudy (anonymous):

@theolympicnomad H always +1 ? i think that's not right.In NaH?

OpenStudy (anonymous):

cibychak you are now splitting hair!!!

OpenStudy (anonymous):

@Kryten okay.But its seriously dangerous to plant a misconception right?

OpenStudy (anonymous):

1) @cibychak any element just by itself not in a compound, their oxidation number is 0 ex: H2's oxidation number is 0 ex: Ag: oxidation number is 0 if its just something like Ag + BLA = LALA 2) the oxidation number of H is always +1, unless its just by itself (see #1) 3) the oxidation number of O is always -2, unless its just by itself (see #1) 4) the oxidation number of F is always -1, unless its just by itself (see#1) ok so after you have written those oxidation numbers in rules 1-4 over each H, F, or O atom in the compound, you can look at the elements that we haven't talked about yet for example:::: N2O4 the oxidation number of O is -2. since there are 4 O's, the charge is -8. now remember that N2O4 has to be neutral so the N2 must have a charge of +8 +8 divided by 2 = +4 N has an oxidation number of +4. more rules: 5) the sum of oxidation numbers in a compound add up to 0 (when multiplied by the subscripts!!!) (see above example) 6) the sum of oxidation numbers in a polyatomic ion is the charge (for example, PO4 has a charge of (-3) so oxidation # of O = -2. (there are 4 O's = -8 charge on that side ) P must have an oxidation number of 5. (-8+5= -3), and -3 is the total charge of the polyatomic ion. THESE ARE VERY SIMPLE RULES THAT YOU HAVE TO REMEMBER. the stuff everyone else is talking about like the group 1 and etc, column 1 or 7, they're just talking about the charges. make sure that you don't listen to them, because other than H, F, and O, most of the elements in their compounds DO NOT have their charge as their oxidation number. in the first example N2O2, N's charge is -3 according to the periodic table, but in the example it turned out to be +4. CHARGES are misleading! And yes it would be dangerous but i am 100% sure this is right :).

OpenStudy (anonymous):

@TheOlympicnomad chill buddy :)

OpenStudy (anonymous):

i can still argue you.just a small correction.the oxidation number of H is always +1, unless its just by itself or when attached to a more electropositive element,like Na :)

OpenStudy (anonymous):

@cibychak lol, :P

OpenStudy (anonymous):

And i still disagree with you. ( ̄ー ̄)